Construct an amortization schedule for a 3-year loan of 50000 at 5% interest, which is to be repaid in quarterly installments over 3 years.
1
Expert's answer
2021-09-07T19:27:31-0400

To get installments we use;

PV"=" PMT"\frac{[1-(1+r)^{-n}]}{r}"

PV"=" 50000

r"=" interest rate"(0.05\/4)=" 0.0125

n "=" the time which is 3

50000 "=" PMT"\frac{[1-(1+0.0125)^{-3}]}{0.0125}"

PMT "=" 17085.06

To get interest for year 1

50000"\times" 0.0125

"=" 625

P"=" installment "-" interest

17085.06"-" 625"=" 16460.06

Outstanding balance is 50000"-" 16460.06

"=" 33539.94

For year 2 we will use first year's balance to calculate interest. Pepe at the same for year 3.